A biquadratic polynomial f(x) has zeroes -2,3,5 and 7. if f(0)=1 then the remainder when f(x) is divided by (x-1)

Grade:7

2 Answers

Arun
4 years ago
Dear student
 
f(x) = a (x+2) (x-3)(x-5)(x-7)
Now f(0) = 1 = – 210 a
a = -1 /210
 
hence f(x) = – (x+2) (x-3)(x-5)(x-7)/210
 
hence by remainder theorem when divisible by (x-1)
remainder = 315/210 = 1.5
Yashraj
3 years ago
since  zeroes are -2,3,5 and 7.
f(x) = a (x+2) (x-3)(x-5)(x-7)
f of o is 1 
so a =-1/210
to find remainder put x=-1
in equation